# Finite Sample Frequency Domain Identification

### v1

@@ Appendix D Proof of Theorem 1
Invoking Lemma 4, we have with probability at least $1-\tilde{\delta}$
$\|V_{k}\|_{\mathrm{op}}$ $\leq\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}+2\tilde{s}\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}$
$\tilde{s}^{2}\mathrm{tr}(\Phi_{v,N}(k))$ $=c^{\prime}\frac{K^{4}}{\sigma_{{e}}^{4}}\|\Phi_{v,N}(k))\|_{\mathrm{op}}(\log(1/\tilde{\delta})+d_{{u}}),$
for some universal constant $c^{\prime}$. Set $\delta/M=\tilde{\delta}$ and consider a union bound over all $\ell\in[M]$. Then, with probability at least $1-\delta$, for all $k=N_{p}\ell$, $\ell\in[M]$
$\|V_{k}\|_{\mathrm{op}}\leq\sqrt{d_{{y}}}\sqrt{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}+\sqrt{c^{\prime}}\frac{K^{2}}{\sigma_{{e}}^{2}}\sqrt{\|\Phi_{v,N}(k))\|_{\mathrm{op}}}\sqrt{\log(M/\delta)+d_{{u}}},$ (25)

### v2

@@ Appendix D Proof of Theorem 1
Invoking Lemma 4, we have
$\mathbb{P}\left(\|V_{k}\|_{\mathrm{op}}\geq(1+2s)\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}\right)\leq 9^{2d_{{u}}}2\exp\left(-\frac{1}{144}s^{2}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right)$
Now fix a failure probability $\tilde{\delta}$. Choosing $s$ such that
$\frac{1}{144}s^{2}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\geq 4d_{{u}}\log 9+\log 1/\tilde{\delta}$
guarantees that the above probability is upper bounded by $\tilde{\delta}$. Hence, with probability at least $1-\tilde{\delta}$
$\|V_{k}\|_{\mathrm{op}}\leq\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}+\sqrt{144}\frac{K^{2}}{\sigma_{{e}}^{2}}\sqrt{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\sqrt{4\log 9\,d_{{u}}+\log 1/\tilde{\delta}}.$
To simplify the final expression define the universal constant
$c=\sqrt{144}\sqrt{4\log 9}.$
Then, we obtain that with probability at least $1-\tilde{\delta}$
$\|V_{k}\|_{\mathrm{op}}\leq\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}+c\frac{K^{2}}{\sigma_{{e}}^{2}}\sqrt{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\sqrt{d_{{u}}+\log 1/\tilde{\delta}}.$ (30)
To finish the proof we need to establish the upper bound (30) over all $k=N_{p}\ell,$ $\ell\in[M]$. It is sufficient to set $\tilde{\delta}=\delta/M$ and consider a union bound over all $\ell\in[M]$. Then, with probability at least $1-\delta$, for all $k=N_{p}\ell$, $\ell\in[M]$
$\|V_{k}\|_{\mathrm{op}}\leq\sqrt{d_{{y}}}\sqrt{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}+c\frac{K^{2}}{\sigma_{{e}}^{2}}\sqrt{\|\Phi_{v,N}(k))\|_{\mathrm{op}}}\sqrt{\log(M/\delta)+d_{{u}}},$